OpenStudy (anonymous):

Jorie is collecting rain water in 30 gallon and 50 gallon barrels. f represents the number of full 30 gallon barrels and g represents the number of full 50 gallon barrels. Which expression describes the total volume of water collected?

OpenStudy (bibby):

You can ask yourself what 10 full 30 gallon barrels means in terms of numbers (30*10=300) Again 10 full 50g barrels = 50*10 = 500

OpenStudy (bibby):

What would the total of 10f + 10g?

OpenStudy (anonymous):

still confused

OpenStudy (bibby):

if f is the number of full 30g barrels than 1f=1*30 2f=2*30 3f=3*30 do you see a pattern?

OpenStudy (bibby):

we set f equal to any number. To get the number of gallons, we multiply f by 30. To get the total gallons in g we multiply by 50 because EACH barrel is 50 gallons.

OpenStudy (anonymous):

ohh thank you